OpenStudy (canada907cat):

If y=cos x, what is y^6 (x)

zepdrix (zepdrix):

You mean \(\large\rm y^{(6)}(x)\) yes? As in 6th derivative?

OpenStudy (canada907cat):

Yes

zepdrix (zepdrix):

Take some derivatives, look for a pattern :D

zepdrix (zepdrix):

Err I should say this... when they give you a similar type problem and ask you to find \(\large\rm y^{(108)}(x)\), then it's imperative that you recognize a pattern, because you don't want to take 108 derivatives.

zepdrix (zepdrix):

But in this case, taking 6 derivatives in a row should be pretty easy.

zepdrix (zepdrix):

Do you remember your sine and cosine derivatives?

OpenStudy (canada907cat):

Well I know that sinx=cos x and cosx=-sinx. Is this what what mean?

zepdrix (zepdrix):

Yes.\[\large\rm \sin x\quad\to\quad \cos x\quad\to\quad-\sin x\quad\to\quad-\cos x\quad\to\quad \sin x\]

OpenStudy (canada907cat):

Is it the same as f(x)=x^(2) which equals 2x?

zepdrix (zepdrix):

I'm not sure what you're asking..

OpenStudy (canada907cat):

Sorry Im looking at something else.

zepdrix (zepdrix):

Anyway, notice that the trig derivatives come full circle after taking 4 of them. See how I started with sine and ended with sine after 4 derivatives?

OpenStudy (canada907cat):

Yes

zepdrix (zepdrix):

So if you take 4 derivatives of cosine, you'll get back to cosine. Which means we really only need to take 2 derivatives. 6 derivatives of cosine is the same as 2 derivatives of cosine.

OpenStudy (canada907cat):

Oh okay so how would that look in terms of the problem?

zepdrix (zepdrix):

\[\large\rm y=\cos x\]\[\large\rm y'=-\sin x\]\[\large\rm y''=-\cos x\]\[\large\rm y'''=\sin x\]\[\large\rm y^{(4)}=\cos x\]So four derivatives got us back to cosine. Take two more derivatives.

OpenStudy (canada907cat):

so y^(5)=-sin x y^(6)=-cox x ?

zepdrix (zepdrix):

Yay good job

OpenStudy (canada907cat):

Wow really! Thank you for explaining that for me.

zepdrix (zepdrix):

np